# Invariant Subspace Decomposition

### v1

@@ 5.2 Time adaptation
In the same setting, we now fix the size of the historical dataset to $n=6000$, which we use to estimate $\hat{\beta}^{\operatorname{inv}}$, and consider an adaptation dataset in which the time-varying coefficients (before the transformation using $U$) undergo two shifts and take values $-0.5$ and $-2$ on two consecutive time windows, each containing $1000$ observations. We assume that the adaptation data are observed sequentially, and use an adaptation window of length $m$ to estimate the residual parameter $\hat{\delta}^{\operatorname{res}}_{t}$ and the OLS solution $\hat{\gamma}^{\operatorname{OLS}}_{t}$. We then compute the squared prediction error of $\hat{\gamma}^{\operatorname{ISD}}_{t}=\hat{\beta}^{\operatorname{inv}}+\hat{\delta}^{\operatorname{res}}_{t}$ and $\hat{\gamma}^{\operatorname{OLS}}_{t}$ on the next data point $X_{t+1}$, i.e., $(X_{t+1}^{\top}(\gamma_{0,t+1}-\hat{\gamma}_{t}))^{2}$ and approximate the corresponding MSPE using a Monte-Carlo approximation with $1000$ draws from $X_{t+1}$. We repeat the simulation $20$ times for different sizes of the adaptation window, $m\in\{1.5p,2p,5p,10p\}$, and plot the obtained MSPE values for $\hat{\gamma}^{\operatorname{OLS}}_{t}$ against $\hat{\gamma}^{\operatorname{ISD}}_{t}$. The result is shown in Figure 6, and empirically supports Theorem 3, in particular that the difference in the MSPEs of the OLS and ISD estimators is larger for small values of $m$ and shrinks for increasing size of the adaptation data (additional details on this simulation are provided in Section A.4 in the Appendix).
Figure 6: MSPE comparison: $\hat{\gamma}^{\operatorname{ISD}}_{t}$ (blue dots) vs. $\hat{\gamma}^{\operatorname{OLS}}_{t}$ for $p=10$ and various adaptation window lengths $m$. The ISD estimator achieves lower MSPE than the OLS for smaller sizes $m$ of the adaptation window, while the two estimators become comparable for increasing $m$. The orange dots show the MSPE of $\beta^{\operatorname{inv}}\text{(oracle)}+\hat{\delta}^{\operatorname{res}}_{t}$ vs. $\hat{\gamma}^{\operatorname{OLS}}_{t}$: if the subspace decomposition is known, then the ISD always achieves lower MSPE than the OLS.
The ISD framework is particularly helpful in scenarios in which the size of the available adaptation window is small (two first plots from the left in Figure 6). A further benefit of the ISD estimator is that it allows us to estimate $\hat{\delta}^{\operatorname{res}}_{t}$ for small lengths $m$ of the adaptation window where $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})<m<p$ and OLS is not feasible.
@@
Figure 7: Average cumulative explained variance on the adaptation data and one standard deviation intervals (over $20$ runs) by the true time-varying parameter $\gamma_{0,t}$ and various estimators. When few data points are available for adaptation at time $t$, the explained variance of the ISD estimator is significantly higher than that of the rolling window OLS, and improves on invariance-based estimators such as the invariant component $\hat{\beta}^{\operatorname{inv}}$ or the maximin $\hat{\beta}^{\operatorname{mm}}$. Due to a shift of $\gamma_{0,t}$ in the residual subspace, the OLS computed on historical data can perform worse than the zero function.
@@ 6 Summary
We propose Invariant Subspace Decomposition (ISD), a framework for invariance-based time adaptation. Our method relies on the orthogonal decomposition of the parameter space into an invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and a residual subspace $\mathcal{S}^{\operatorname{res}}$, such that the maximizer of the explained variance over $\mathcal{S}^{\operatorname{inv}}$ is time-invariant. The estimation of the invariant component $\beta^{\operatorname{inv}}$ on a large historical dataset and the reduced dimensionality of $\mathcal{S}^{\operatorname{res}}$ with respect to the original parameter space $\mathbb{R}^{p}$ allow the ISD estimator to improve on the prediction accuracy of existing estimation techniques. We provide finite sample guarantees for the proposed estimation method and additionally support the validity of our theoretical results through simulated experiments.
@@ Acknowledgments
NP and ML are supported by a research grant (0069071) from Novo Nordisk Fonden.
@@ A.1 Example of non-uniqueness of an irreducible orthogonal partition
Define for all $j\in\{1,2,3\}$ the linear spaces $\mathcal{S}_{j}=\langle e_{j}\rangle$, where $e_{j}$ is the $j$-th vector of the canonical basis. Then since $\Sigma^{1}$ and $\Sigma^{2}$ are (block) diagonal the partition $\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}$ is an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. Consider now the orthonormal matrix
@@
also form an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ (this follows, for example, from Proposition 1) but $\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}\neq\{\tilde{\mathcal{S}}_{1},\tilde{\mathcal{S}}_{2},\tilde{\mathcal{S}}_{3}\}$.
@@
A.2 Threshold selection for opt-invariant subspaces
where $\hat{c}_{k}^{j}\coloneqq\widehat{\operatorname{Corr}}(\mathbf{Y}_{k}-\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}),\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}_{k}))$. We then split the historical data into $J=10$ disjoint blocks of observations, and for all $j\in\{1,\dots,J\}$ denote by $\mathbf{X}_{j}$ and $\mathbf{Y}_{j}$ the observations in the $j$-th block and by $\mathbf{X}_{-j}$ and $\mathbf{Y}_{-j}$ the remaining historical data. For all possible thresholds $\lambda\in\Lambda$ we proceed in the following way. For all folds $j\in\{1,\dots,J\}$, we compute an estimate for the invariant component as in (19) using $\mathbf{X}_{-j}$ and $\mathbf{Y}_{-j}$, which we denote by $\hat{\beta}^{\operatorname{inv},-j}(\lambda)$. Inside the left-out $j$-th block of observations, we then consider a rolling window of length $d=2p$ and the observation at $t^{*}$ immediately following the rolling window: we compute the residual parameter $\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)$ as in (21) using the $d$ observations in the rolling window, and evaluate the empirical explained variance by $\hat{\beta}^{\operatorname{inv},-j}(\lambda)+\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)$ on the observation at $t^{*}$, i.e.,
$\widehat{\operatorname{\Delta Var}}_{t^{*}}(\lambda)\coloneqq(Y_{t^{*}})^{2}-(Y_{t^{*}}-X_{t^{*}}^{\top}(\hat{\beta}^{\operatorname{inv},-j}(\lambda)+\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)))^{2}.$
We repeat this computation for all possible $t^{*}\in\mathcal{I}_{j}$, where $\mathcal{I}_{j}$ denotes the time points in the $j$-th block of observations excluding the first $d$ observations, and define $\overline{\operatorname{\Delta Var}}_{j}^{\lambda}\coloneqq\frac{1}{|\mathcal{I}_{j}|}\sum_{t^{*}\in\mathcal{I}_{j}}\widehat{\operatorname{\Delta Var}}_{t^{*}}(\lambda)$. For all $\lambda\in\Lambda$, we denote the average explained variance over the $J$ folds by $\overline{\operatorname{\Delta Var}}^{\lambda}\coloneqq\frac{1}{J}\sum_{j=1}^{J}\overline{\operatorname{\Delta Var}}_{j}^{\lambda}$ and the standard error (across the $J$ folds) of such explained variance as $\operatorname{se}(\overline{\operatorname{\Delta Var}}^{\lambda})\coloneqq\frac{1}{J}\sqrt{\sum_{j=1}^{J}(\overline{\operatorname{\Delta Var}}_{j}^{\lambda}-\overline{\operatorname{\Delta Var}}^{\lambda})^{2}}$. Moreover, let $\lambda^{\max}\coloneqq\operatorname*{arg\,max}_{\lambda\in\Lambda}\overline{\operatorname{\Delta Var}}^{\lambda}$. Then, we choose the optimal threshold as
@@ A.3 Methods and computational complexity of joint block diagonalization
However, in general, methods for JBD cannot be directly applied to solve AJBD. Algorithms that solve AJBD directly – based on the iterative optimization of a cost function via matrix rotations – have been developed, for example, by Tichavsky and Koldovsky (2012) and Févotte and Theis (2007), but require the number of diagonal blocks to be known in advance. Alternatively and similarly to how JBD can be solved by AJD methods, one can also, with some slight modifications, use AJD methods to solve AJBD. More specifically, some heuristics need to be used to determine the size of the blocks: these can consist, for example, in setting a threshold for the non-zero off-block-diagonal elements in the transformed matrices.
@@
A.4 Further simulation details: MSPE comparison

### v2

@@ 5.2 Time adaptation
In the same setting, we now fix the size of the historical dataset to $n=6000$, which we use to estimate $\hat{\beta}^{\operatorname{inv}}$, and consider a test dataset in which the time-varying coefficients (before the transformation using $U$) undergo two shifts and take values $-0.5$ and $-2$ on two consecutive time windows, each containing $1000$ observations. We assume that the test data are observed sequentially, and take as adaptation data a rolling window of length $m$ contained in the test data and shifting by one time point at the time. We use these sequential adaptation datasets to estimate the residual parameter $\hat{\delta}^{\operatorname{res}}_{t}$ and the OLS solution $\hat{\gamma}^{\operatorname{OLS}}_{t}$. We then compute the squared prediction error of $\hat{\gamma}^{\operatorname{ISD}}_{t}=\hat{\beta}^{\operatorname{inv}}+\hat{\delta}^{\operatorname{res}}_{t}$ and $\hat{\gamma}^{\operatorname{OLS}}_{t}$ on the next data point $X_{t+1}$, i.e., $(X_{t+1}^{\top}(\gamma_{0,t+1}-\hat{\gamma}_{t}))^{2}$ and approximate the corresponding MSPE using a Monte-Carlo approximation with $1000$ draws from $X_{t+1}$ (these correspond to the $1000$ sequential observations in each of the two windows of the test data). We repeat the simulation $20$ times for different sizes of the adaptation window, $m\in\{1.5p,2p,5p,10p\}$, and plot the obtained MSPE values for $\hat{\gamma}^{\operatorname{OLS}}_{t}$ against $\hat{\gamma}^{\operatorname{ISD}}_{t}$. The result is shown in Figure 6, and empirically supports Theorem 3, in particular that the difference in the MSPEs of the OLS and ISD estimators is proportional to the ratio $\frac{\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}{m}$ and is therefore larger for small values of $m$ and shrinks for increasing size of the adaptation data (additional details on this simulation are provided in Section A.5 in the Appendix).
Figure 6: MSPE comparison: $\hat{\gamma}^{\operatorname{ISD}}_{t}$ (blue dots) vs. $\hat{\gamma}^{\operatorname{OLS}}_{t}$ for $p=10$ and various adaptation window lengths $m$ (see Section 5.2). The ISD estimator achieves lower MSPE than the OLS for smaller sizes $m$ of the adaptation window, while the two estimators become comparable for increasing $m$. The orange dots show the MSPE of $\beta^{\operatorname{inv}}\text{(oracle)}+\hat{\delta}^{\operatorname{res}}_{t}$ vs. $\hat{\gamma}^{\operatorname{OLS}}_{t}$: if the subspace decomposition is known, then the ISD always achieves lower MSPE than the OLS.
The ISD framework is particularly helpful in scenarios in which the size of the available adaptation window is small (two first plots from the left in Figure 6). Indeed, from Theorem 3 it also follows that the larger the dimension of the invariant subspace the greater the advantage in using the ISD framework for prediction rather than a naive OLS approach. A further benefit of the ISD estimator is that it allows us to estimate $\hat{\delta}^{\operatorname{res}}_{t}$ for small lengths $m$ of the adaptation window where $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})<m<p$ and OLS is not feasible.
@@
Figure 7: Average cumulative explained variance on the adaptation data and one standard deviation intervals (over $20$ runs) by the true time-varying parameter $\gamma_{0,t}$ and various estimators (see Section 5.2). When few data points are available for adaptation at time $t$, the explained variance of the ISD estimator is significantly higher than that of the rolling window OLS, and improves on invariance-based estimators such as the invariant component $\hat{\beta}^{\operatorname{inv}}$ or the maximin $\hat{\beta}^{\operatorname{mm}}$. Due to a shift of $\gamma_{0,t}$ in the residual subspace, the OLS computed on historical data can perform worse than the zero function.
5.3 Real data example
We now present a toy example that applies ISD to real data. The data used for this experiment are collected using a controlled physical system developed by Gamella et al. (2025). The system, shown in Figure 8, consists of a light tunnel with a light source $X_{\operatorname{RGB}}$ whose emitted light passes through two polarizers with relative angle $\theta$ between them and is captured by a sensor placed at the end of the tunnel, behind the polarizers. At the end of the tunnel there are two additional LED light sources, $X_{L_{31}}$ and $X_{L_{32}}$, whose emitted light is unaffected by the polarizers. The sensor $Y:=\tilde{I}_{3}$ measures the overall infrared light at the end of the tunnel, which is affected by the intensity of the RGB source and by the two LEDs. As described by Gamella et al. (2025), the effect of $X_{\operatorname{RGB}}$ on $Y$ is linear and depends on $\theta$, more precisely, it holds that $Y\propto\operatorname{cos}^{2}(\theta)X_{\operatorname{RGB}}$ The dependence of $Y$ on the two LEDs is instead independent of $\theta$. We take $Y$ as our response, consider the covariates vector $[X_{\operatorname{RGB}},X_{L_{31}},X_{L_{32}}]^{\top}\in\mathbb{R}^{3}$ and assume that the angle $\theta$ is unknown. Since we control the three light sources independently, and we expect the dependence of the infrared light $Y$ on the two LEDs to remain the same across time, we hope to detect a nontrivial invariant subspace related to the two LED covariates.
Figure 8: Illustration of the light tunnel, see Section 5.3. The figure is taken from Gamella et al. (2025) (published under a Creative Commons Attribution 4.0 International License (https://creativecommons.org/licenses/by/4.0/)). The variables of interest in our experiment are the $RGB$ values of the light source, the LEDs intensities $L_{31}$ and $L_{32}$, and the measurement of the light sensor at the end of the tunnel $\tilde{I}_{3}$.
The available dataset contains $8000$ observations, collected under changing values of the angle $\theta$. The historical dataset contains the first $7000$ observations, and the test dataset the remaining $1000$. Figure 9 shows the dependence of the response on the three covariates, as well as the values of the response and of $\operatorname{cos}^{2}(\theta)$ through time.
Figure 9: Dataset of the experiment discussed in Section 5.3. The top four figures from the left show the dependence of the infrared measurement $Y$ on the covariates and on time $t$ (encoded by the colormap) for historical data. The angle of the polarizers changes over time (top right) and thus has an influence on the linear relationship between $Y$ and $X_{\operatorname{RGB}}$ (top left). For this experiment we assume that the angle $\theta$ is unknown. The second row shows the same quantities during test time, where the polarization angle is much smaller (bottom right). The dependence between $Y$ and the two LEDs is small but significant (testing a reduced linear model without either $L_{31}$ or $L_{32}$ against the full model results in $p$-values smaller than $10^{-4}$, both for historical and test data).
We apply ISD on the historical data to find an invariant component $\hat{\beta}^{\operatorname{inv}}$. For comparison, we also compute the maximin $\hat{\beta}^{\operatorname{mm}}$(Bühlmann and Meinshausen, 2015) and the OLS solution $\hat{\beta}^{\operatorname{OLS}}$ on historical data. We then use these estimated parameters for zero-shot prediction on test data. Table 1 shows the $R^{2}$ coefficient, defined as in Section 5.1 as the fraction of explained variance $R^{2}=\frac{\sum_{t=1}^{n}(\widehat{\operatorname{Var}}(Y_{t})-\widehat{\operatorname{Var}}(Y_{t}-X_{t}^{\top}\hat{\beta}))}{\sum_{t=1}^{n}\widehat{\operatorname{Var}}(Y_{t})}$, on historical and test data.
$R^{2}$ historical data test data (zero-shot prediction)
$\hat{\beta}^{\operatorname{inv}}$ 0.019 0.062
$\hat{\beta}^{\operatorname{OLS}}$ 0.558 -10.703
$\hat{\beta}^{\operatorname{mm}}$ 0.477 -2.194
Table 1: Normalized explained variance ($R^{2}$) by $\hat{\beta}^{\operatorname{inv}}$ and comparison with $\hat{\beta}^{\operatorname{OLS}}$ and $\hat{\beta}^{\operatorname{mm}}$: training (historical data) and zero-shot generalization (test data); see Section 5.3.
The invariant component $\hat{\beta}^{\operatorname{inv}}$ is the only estimator that achieves positive explained variance on test data. This is because $\hat{\beta}^{\operatorname{inv}}$ only captures the parts of the variance that can be transferred to the test data, as can be seen from the lower explained variance compared to $\hat{\beta}^{\operatorname{OLS}}$ on the historical data. The estimated invariant subspace $\hat{\mathcal{S}}^{\operatorname{inv}}=\operatorname{span}\{[-0.13910168,-0.95836843,-0.24936055]^{\top}\}$ has dimension $1$ and shows in particular that most of the invariant information is encoded in the two LEDs, with highest weight given to $X_{L_{31}}$. This is expected since the LEDs intensities are not affected by the changing angle between the two polarizers. Also the relatively small $R^{2}$ is expected: since the RGB light source is stronger than the two LEDs at the end of the tunnel, most of the variance in $Y$ is explained by the non-invariant component $X_{\operatorname{RGB}}$ (see Figure 9). However, it is not the case that the whole subspace spanned by the two covariates corresponding to the LEDs is invariant, as we would have assumed from the knowledge of the physical system. An explanation is that due to the data collection process there is nonzero observed correlation between $X_{\operatorname{RGB}}$ and $X_{L_{32}}$, but not between $X_{\operatorname{RGB}}$ and $X_{L_{31}}$ (on the historical data, we have $\operatorname{corr(X_{\operatorname{RGB}},X_{L_{31}})}=0.004$ with p-value $0.723$ and $\operatorname{corr(X_{\operatorname{RGB}},X_{L_{32}})}=-0.107$ with p-value smaller than $10^{-15}$).
We also run the adaptation step considering as adaptation data a rolling window of size $m=8$ shifting through the test data. We show in Figure 10 the cumulative explained variance obtained by ISD, by the OLS solution computed on the same adaptation data and by the OLS solution and the invariant component computed on historical data. The plot shows that $\hat{\gamma}^{ISD{}}$ achieves the highest explained variance, with a small improvement on the rolling window OLS $\hat{\gamma}^{OLS}$. Indeed, in this particular example the size of the invariant subspace is small ($\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})=1$) and by Theorem 3, we expect only a small improvement.
Figure 10: Cumulative explained variance on the adaptation data by the time adaptation estimators $\hat{\gamma}^{\operatorname{ISD}}$ and $\hat{\gamma}^{\operatorname{OLS}}$ and by the zero-shot predictors $\hat{\beta}^{\operatorname{inv}}$ and $\hat{\beta}^{\operatorname{OLS}}$; see Section 5.3.
@@ 6 Summary
We propose Invariant Subspace Decomposition (ISD), a framework for invariance-based time adaptation. Our method relies on the orthogonal decomposition of the parameter space into an invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and a residual subspace $\mathcal{S}^{\operatorname{res}}$, such that the maximizer of the explained variance over $\mathcal{S}^{\operatorname{inv}}$ is time-invariant. The estimation of the invariant component $\beta^{\operatorname{inv}}$ on a large historical dataset and the reduced dimensionality of $\mathcal{S}^{\operatorname{res}}$ with respect to the original parameter space $\mathbb{R}^{p}$ allow the ISD estimator to improve on the prediction accuracy of existing estimation techniques. We provide finite sample guarantees for the proposed estimation method and additionally support the validity of our theoretical results through simulated experiments and one real world data experiment.
@@ Acknowledgments
We thank Juan Gamella for generating the data for the presented real world example, and the three anonymous reviewers for the valuable comments. NP and ML are supported by a research grant (0069071) from Novo Nordisk Fonden.
@@ A.1 Example of non-uniqueness of an irreducible orthogonal partition
Define for all $j\in\{1,2,3\}$ the linear spaces $\mathcal{S}_{j}=\langle e_{j}\rangle$, where $e_{j}$ is the $j$-th vector of the canonical basis. Then since $\Sigma^{1}$ and $\Sigma^{2}$ are (block) diagonal the partition $\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}$ is an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Consider now the orthonormal matrix
@@
also form an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition (this follows, for example, from Proposition 1) but $\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}\neq\{\tilde{\mathcal{S}}_{1},\tilde{\mathcal{S}}_{2},\tilde{\mathcal{S}}_{3}\}$.
@@
A.2 Example of quickly varying $\gamma_{0,t}$ and zero-shot generalization
We repeat the same simulation described in Section 5.1 but now consider non-smooth variations of $\gamma_{0,t}$. More specifically, the only difference from the simulation described in Section 5.1 is that we let the $3$ time-varying entries of $\gamma_{0,t}$ (prior to its rotation by $U$) vary quickly with $t$.
Figure 12: Normalized explained variance ($R^{2}$) by $\hat{\beta}^{\operatorname{inv}}$ and comparison with $\beta^{\operatorname{inv}}$, $\hat{\beta}^{\operatorname{mm}}$ and $\hat{\beta}^{\operatorname{OLS}}$: training (historical data, left) and zero-shot generalization (test data, right), for different sizes $n$ of the historical data. The dashed line indicates the population value of the (normalized) explained variance by $\beta^{\operatorname{inv}}$. In the historical data, $\gamma_{0,t}$ is quickly varying.
More specifically, at each time point each one of the $3$ entries is sampled uniformly in an interval of width $1$ centered around a value which changes $20$ times in the observed time horizon $n$. These centers are randomly sampled in $[0,1.2]$. These choices of the sampling intervals ensure that the experiments are comparable for different sizes $n$ of the historical data and that there is a shift in $\gamma_{0,t}$ outside of the observed support in the test data (which is generated as in Section 5.1). This experiment supports Remark 2, showing that we do not need assumptions on the type of changes in $\gamma_{0,t}$ in the historical data.
In particular, Figure 11 shows that the MSE of $\hat{\beta}^{\operatorname{inv}}$ converges to zero for increasing values of $n$, that is, we are able to estimate the invariant component even when $\gamma_{0,t}$ is quickly varying in the historical data.
Moreover, Figure 12 shows the results in terms of the $R^{2}$ coefficient. In particular, for $n$ large enough, the $R^{2}$ coefficient of the estimated invariant component $\hat{\beta}^{\operatorname{inv}}$ remains positive even for values of the test data that lie outside of the observed support.
@@ A.3 Methods and computational complexity of joint block diagonalization
However, in general, methods for JBD cannot be directly applied to solve AJBD. Algorithms that solve AJBD directly—based on the iterative optimization of a cost function via matrix rotations—have been developed, for example, by Tichavsky and Koldovsky (2012) and Févotte and Theis (2007), but require the number of diagonal blocks to be known in advance. Alternatively and similarly to how JBD can be solved by AJD methods, one can also, with some slight modifications, use AJD methods to solve AJBD. More specifically, some heuristics need to be used to determine the size of the blocks: these can consist, for example, in setting a threshold for the non-zero off-block-diagonal elements in the transformed matrices.
@@
A.4 Threshold selection for opt-invariant subspaces
where, for all $j\in\{1,\dots,q^{\hat{U}}_{\max}\}$ and $k\in\{1,\dots,K\}$, $\hat{c}_{k}^{j}\coloneqq\widehat{\operatorname{Corr}}(\mathbf{Y}_{k}-\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}),\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}))$. We then split the historical data into $L=10$ disjoint blocks of observations, and for all $j\in\{1,\dots,J\}$ denote by $\mathbf{X}_{\ell}$ and $\mathbf{Y}_{\ell}$ the observations in the $\ell$-th block and by $\mathbf{X}_{-\ell}$ and $\mathbf{Y}_{-\ell}$ the remaining historical data. For all possible thresholds $\lambda\in\Lambda$ we proceed in the following way. For all folds $\ell\in\{1,\dots,L\}$, we compute an estimate for the invariant component as in (20) using $\mathbf{X}_{-\ell}$ and $\mathbf{Y}_{-\ell}$, which we denote by $\hat{\beta}^{\operatorname{inv},-\ell}(\lambda)$. Inside the left-out $\ell$-th block of observations, we then consider a rolling window of length $d=2p$ and the observation at $t^{*}$ immediately following the rolling window: we compute the residual parameter $\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)$ as in (22) using the $d$ observations in the rolling window, and evaluate the empirical explained variance by $\hat{\beta}^{\operatorname{inv},-\ell}(\lambda)+\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)$ on the observation at $t^{*}$, i.e.,
$\widehat{\operatorname{\Delta Var}}_{t^{*}}(\lambda)\coloneqq(Y_{t^{*}})^{2}-(Y_{t^{*}}-X_{t^{*}}^{\top}(\hat{\beta}^{\operatorname{inv},-\ell}(\lambda)+\hat{\delta}^{\operatorname{res}}_{t^{*}}(\lambda)))^{2}.$
We repeat this computation for all possible $t^{*}\in\mathcal{I}_{\ell}$, where $\mathcal{I}_{\ell}$ denotes the time points in the $\ell$-th block of observations excluding the first $d$ observations, and define $\overline{\operatorname{\Delta Var}}_{\ell}^{\lambda}\coloneqq\frac{1}{|\mathcal{I}_{\ell}|}\sum_{t^{*}\in\mathcal{I}_{\ell}}\widehat{\operatorname{\Delta Var}}_{t^{*}}(\lambda)$. For all $\lambda\in\Lambda$, we denote the average explained variance over the $L$ folds by $\overline{\operatorname{\Delta Var}}^{\lambda}\coloneqq\frac{1}{L}\sum_{\ell=1}^{L}\overline{\operatorname{\Delta Var}}_{\ell}^{\lambda}$ and the standard error (across the $L$ folds) of such explained variance as $\operatorname{se}(\overline{\operatorname{\Delta Var}}^{\lambda})\coloneqq\frac{1}{L}\sqrt{\sum_{\ell=1}^{L}(\overline{\operatorname{\Delta Var}}_{\ell}^{\lambda}-\overline{\operatorname{\Delta Var}}^{\lambda})^{2}}$. Moreover, let $\lambda^{\max}\coloneqq\operatorname*{arg\,max}_{\lambda\in\Lambda}\overline{\operatorname{\Delta Var}}^{\lambda}$. Then, we choose the optimal threshold as
A.5 Further simulation details: MSPE comparison